# Tradeoffs of Diagonal Fisher Information Matrix Estimators

### v2

@@ 4 Practical Variance Estimation
The neural network mapping w.r.t. the $j$’th neuron in the second last layer is $\bm{h}(\bm{x})=\bm{w}_{-1}^{j}\phi(\bm{h}^{\top}_{-2}(\bm{x})\bm{w}_{-2}^{j}+C_{-2})+\bm{c}_{-1}$, where $\bm{w}_{-2}^{j}$ and $\bm{w}_{-1}^{j}$ are incoming and outgoing links of the interested neuron, respectively; $\bm{h}_{-2}(\bm{x})$ is the output of the third last layer; and $\phi$ is the activation function. The ‘constants’ $C_{-2}$ and $\bm{c}_{-1}$ denote an aggregation of all terms which are independent of $\bm{w}_{-2}^{j}$ and $\bm{w}_{-1}^{j}$ in their respective layers. The diagonal Hessian w.r.t. $\bm{w}_{-2}^{j}$ is $\partial^{2}\bm{h}(\bm{x})=\bm{w}_{-1}^{j}\phi^{\prime\prime}(\bm{h}^{\top}_{-2}(\bm{x})\bm{w}_{-2}^{j}+C_{-2})(\bm{h}_{-2}(\bm{x})\circ\bm{h}_{-2}(\bm{x}))^{\top}$. By Theorem 4.1, ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ can be arbitrarily small depending on $\phi^{\prime\prime}(\bm{h}^{\top}_{-2}(\bm{x})\bm{w}_{-2}^{j}+C_{-2})$. For example, if $\phi(t)=1/(1+\exp(-t))$, then $\phi^{\prime\prime}(t)=\phi(t)(1-\phi(t))(1-2\phi(t))$. In this case, for a neuron in the second last layer, a sufficient condition for ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})=0$ (and having ^$\hat{\mathcal{I}}_{2}$ favored against ^$\hat{\mathcal{I}}_{1}$) is for the neuron’s pre-activation value $\bm{h}^{\top}_{-2}(\bm{x})\bm{w}_{-2}^{j}+C_{-2}=0$. When the pre-activation value is saturated ($\infty$- ∞ / $\infty$), we also have that ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})={\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})=0$. Alternatively, suppose that $\phi(t)=\textrm{SoftPlus}(t)\vcentcolon=\log(1+\exp(t))$, a continuous relaxation of $\mathrm{ReLU}$, then $\phi^{\prime\prime}(t)=\phi^{\prime}(t)(1-\phi^{\prime}(t))$ where $\phi^{\prime}(t)=1/(1+\exp(-t))$. As such, for a neuron in the second last layer, a sufficient condition for ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})=0$ with ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})\neq 0$, is when $\bm{h}^{\top}_{-2}(\bm{x})\bm{w}_{-2}^{j}+C_{-2}\rightarrow\infty$.
These observations can be further clarified by taking a look at these quantities over multiple parameters. So far we have only examined the variance of the FIM element-wise w.r.t. to the parameters of $\bm{h}$. To study all parameters jointly, we consider the trace variances of the FIM estimators: for any $j\in\{1,2\}$, $\mathcal{V}_{j}(\bm{\theta}\,|\,\bm{x})$ denotes the trace of the covariance matrix of $\operatorname*{\mathrm{diag}}(\hat{\mathcal{I}}_{j}(\bm{\theta}\,|\,\bm{x}))$, where $\operatorname*{\mathrm{diag}}(\cdot)$ extracts the diagonal elements of the matrix into a column vector. In what follows, we present upper bounds of these joint quantities.
@@
where $\|\cdot\|_{F}$ denotes the Frobenius norm and $\mathrm{dHes}(\bm{h}\,|\,\bm{x})=(\operatorname*{\mathrm{diag}}(\mathrm{Hes}(\bm{h}_{1}\,|\,\bm{x})),\ldots,\operatorname*{\mathrm{diag}}(\mathrm{Hes}(\bm{h}_{T}\,|\,\bm{x})))$.
This upper bound comes from a combination of simply integrating the parameter-wise variances in Theorem 4.1 and incorporating an alternative trace variance bound which takes into account the full spectrum of the network derivatives and sufficient statistics quantities, full depicted in Theorem C.1. Lower bounds can also be derived in terms of singular values, we leave these to the Appendix for conciseness. We also note that the upper bound shown in Corollary 4.6 can also be improved by expressing the first term of the $\min$ function as singular value quantities.
Having the $\min$ function in Corollary 4.6 can be helpful as it shows a trade-off in two different upper bounds: the scale of the network derivatives $\partial\bm{h}(\bm{x})$ and $\partial^{2}\bm{h}(\bm{x})$ with the spectrum of the sufficient statistics terms. In the case of Eqs. 17 and 19, the trace of the $\mathcal{I}(\bm{h}\,|\,\bm{x})$ is exactly the sum of all eigenvalues, including $\lambda_{\max}(\mathcal{I}(\bm{h}\,|\,\bm{x}))$. This can be helpful as an alternative bound when the scale of the network derivatives are not guaranteed to be bounded by a small value. It should be noted, that by the chain rule, the scale of these network derivatives directly scale with the overall sharpness / flatness (Li et al., 2018) of the network which is associated with the derivatives of the loss function, i.e., the log-likelihood of Eq. 1. For networks with large network derivative, the first term of the $\min$ is preferred to prevent quadratic scaling of $\|\partial\bm{h}(\bm{x})\|$. On the other hand, if the sharpness of the network $\bm{h}$ can be controlled, e.g. via sharpness aware minimization (Foret et al., 2020), then the second term of the $\min$ might be preferred to avoid considering the full spectrum of $\mathcal{I}(\bm{h}\,|\,\bm{x})$ in the bound.
4.1 Joint FIM Estimators
In the above, we considered the variance of conditional FIMs, which can scale differently depending on the input $\bm{x}$. Nevertheless, the ‘joint FIM’ estimators ^$\hat{\mathcal{I}}_{j}(\theta_{i})$ depend on sampling of $\bm{x}$ w.r.t. the data distribution $q(\bm{x})$. The bounds presented in Eq. 10 can be extended to the joint FIM $\mathcal{I}(\theta_{i})\vcentcolon=\mathop{\mathbb{E}}_{q(\bm{x})}\mathcal{I}(\theta_{i}\,|\,\bm{x})$ by simply taking an expectation $\mathop{\mathbb{E}}_{q(\bm{x})}$ uniformly over the bounds. To analysis the variances of the joint FIM estimators ${\mathcal{V}}_{j}(\theta_{i})$, we present the following theorem which connects the prior results established for ${\mathcal{V}}_{j}(\theta_{i}\,|\,\bm{x})$, e.g. Theorem 4.1, via the law of total variance.
@@ 4.1 Joint FIM Estimators
For $j\in\{1,2\}$,
${\mathcal{V}}_{j}(\theta_{i})=\frac{1}{N}\cdot\mathrm{Var}\left(\mathcal{I}(\theta_{i}\,|\,\bm{x})\right)+\mathop{\mathbb{E}}_{q(\bm{x})}\left[{\mathcal{V}}_{j}(\theta_{i}\,|\,\bm{x})\right],$ (20)
where $\mathrm{Var}\left(\mathcal{I}(\theta_{i}\,|\,\bm{x})\right)$ is the variance of $\mathcal{I}(\theta_{i}\,|\,\bm{x})$ w.r.t. $q(\bm{x})$ and $N$ is the number of samples of $\bm{y}\sim p(\bm{y}\,|\,\bm{x})$ for each $\bm{x}$.
The variance incurred when estimating the FIM has two components. The first term on the RHS of Eq. 20 characterizes the randomness of the FIM w.r.t. $q(\bm{x})$, or the input randomness. It vanishes when the FIM is estimated by taking the expectation w.r.t. $q(\bm{x})$, or the number of samples $\bm{x}_{k}$ is large enough. The second term is caused by the sampling of $\bm{y}_{k}$ according to $p(\bm{y}\,|\,\bm{x})$, or the output randomness, which scales with the central moments of $\bm{t}(\bm{y})$. If the network is trained so that $p(\bm{y}\,|\,\bm{x})$ tends to be deterministic, this term will disappear leaving the first term to dominate.
Eq. 20 can be further generalized using the law of total covariance to extend prior work considering conditional FIM covariances Soen & Sun (2021) to joint FIM covariances.
Notably, Theorem 4.7 connects the variance of assuming a fixed input $\bm{x}$ with multiple samples $\bm{y}_{k}$ with the variance of pairs of samples $(\bm{x}_{k},\bm{y}_{k})$. The bounds for ${\mathcal{V}}_{j}(\theta_{i}\,|\,\bm{x})$ explored in this section can thus be applied to the corresponding joint variance ${\mathcal{V}}_{j}(\theta_{i})$ by using this theorem. This is straightforward and omitted.
The first variance term of Theorem 4.7 cannot be computed in practice: it relies on how the closed-form FIM varies w.r.t. the input distribution $q(\bm{x})$. As such, it is useful to bound the first term into computable quantities.
@@
One may notice that this upper bound is very similar to the 4th central moment term $\lambda_{\max}(\mathcal{K}^{p}(\bm{t}\,|\,\bm{x}))$ when considering the upper bound of estimator variance ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ when applying Theorems 4.1 and 4.2. As such, one may wonder when $\lambda^{2}_{\max}(\mathcal{I}(\bm{h}\,|\,\bm{x}))$ and $\mathcal{K}^{p}_{\max}(\bm{t}\,|\,\bm{x})$ are equivalent. The equality $\lambda^{2}_{\max}(\mathcal{I}(\bm{h}\,|\,\bm{x}))=\mathcal{K}^{p}_{\max}(\bm{t}\,|\,\bm{x})$ holds iff $p(\bm{y}\,|\,\bm{x})$ does not depend on the outputs of $\bm{h}(\bm{x})$. Thus in our case, where a neural network parameterizes our exponential family, theses eigenvalues will be distinct.
In terms of sample complexity of the joint variance, the complexity scales inversely to the joint number of samples. For instance, assuming that we take empirical expectations with $N_{x}$ samples (or that $q(\bm{x})$ consists of the sum of Dirac delta functions) and each $\bm{x}$ generates $N_{y}$ samples of $\bm{y}\sim p(\bm{y}\,|\,\bm{x})$, using Theorems 4.7 and 4.8, then the bound sample complexity is $\mathcal{O}(1/(N_{x}N_{y}))$.
@@ 5 Case Studies
To make the implications of our theoretic results more concrete, we consider two common settings in supervised learning: regression and classification, which corresponds respectively to specifying the exponential family distribution in Eq. 1 as a multi-dimensional isotropic Gaussian distribution or a categorical distribution. We also show empirical evidence drawn from an MNIST classifier network.
5.1 Regression: Isotropic Gaussian Distribution
Consider the family of multi-dimensional isotropic Gaussian distribution. As per Eq. 1, we have $\bm{t}(\bm{y})=\bm{y}\in\Re^{T}$ and base measure $\pi(\bm{y})\propto\exp(-\frac{1}{2}{\bm{y}}^{\top}{\bm{y}})$. This corresponds to the case where $\bm{h}(\bm{x})$ is trained via the squared loss. The corresponding central moments of $\bm{t}(\bm{y})$ are:
$\mathcal{I}(\bm{h}\,|\,\bm{x})$ $=\bm{I};\;\textrm{and}\;$
$\mathcal{K}_{abcd}^{p}(\bm{t}\,|\,\bm{x})$ $=\mathcal{I}_{ab}(\bm{h}\,|\,\bm{x})\cdot\mathcal{I}_{cd}(\bm{h}\,|\,\bm{x})+\,\mathcal{I}_{ac}(\bm{h}\,|\,\bm{x})\cdot\mathcal{I}_{bd}(\bm{h}\,|\,\bm{x})+\mathcal{I}_{ad}(\bm{h}\,|\,\bm{x})\cdot\mathcal{I}_{bc}(\bm{h}\,|\,\bm{x}).$
The derivatives of the corresponding log-partition function $F(\bm{h})$ yields these central moments (Soen & Sun, 2021, Lemma 5). To apply Theorem 4.1, we examine the extreme eigenvalues of $\mathcal{I}(\bm{h}\,|\,\bm{x})$ and $\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})-\mathcal{I}(\bm{h}\,|\,\bm{x})\otimes\mathcal{I}(\bm{h}\,|\,\bm{x})$.
@@ 5.1 Regression: Isotropic Gaussian Distribution
Suppose that the exponential family in Eq. 1 is an isotropic Gaussian distribution. Then:
@@
$\tilde{\lambda}_{\min}(\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})-\mathcal{I}(\bm{h}\,|\,\bm{x})\otimes\mathcal{I}(\bm{h}\,|\,\bm{x}))$ $=2;$
$\tilde{\lambda}_{\max}(\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})-\mathcal{I}(\bm{h}\,|\,\bm{x})\otimes\mathcal{I}(\bm{h}\,|\,\bm{x}))$ $={2}.$
@@
In the case of an isotropic Gaussian distribution, the eigenvalues of both sufficient statistics quantities for our bounds are equal. As such the bound for $\mathcal{I}(\theta_{i}\,|\,\bm{x})$, ^$\hat{\mathcal{I}}_{1}(\bm{h}\,|\,\bm{x})$, and ^$\hat{\mathcal{I}}_{2}(\bm{h}\,|\,\bm{x})$ in Theorem 4.1 are all equalities for standard regression.
@@
As $\mathcal{I}(\theta_{i}\,|\,\bm{x})$ can be written exactly in terms of the gradients of $\bm{h}$, in practice one does not need to utilize a random estimator when computing the conditional FIM. Indeed, the conditional FIM simplifies to a Gauss-Newton matrix for the squared loss (Martens, 2020). When computing the FIM over a random sample $\bm{x}$, a variance still appears due to Theorem 4.7. Utilizing Lemma 4.8 and Proposition 5.1, the variance over a joint distribution is bounded by a function of the derivative $\partial_{i}\bm{h}(\bm{x})$ over the marginal input distributions:
${\mathcal{V}}(\theta_{i})\leq\mathop{\mathbb{E}}_{q(\bm{x})}[\|\partial_{i}\bm{h}(\bm{x})\|_{2}^{4}]\leq\max_{\bm{x}\in\mathrm{Supp}(q)}\|\partial_{i}\bm{h}(\bm{x})\|_{2}^{4}.$
In other words, the overall variance to approximate the joint FIM is bounded by the gradients of the network outputs.
5.2 Classification: Categorical Distribution
@@
For multi-class classification, we instantiate our exponential family with a categorical distribution (with unit base measure $\pi(y)=1$). This corresponds to training a classifier network with log-loss. Let $\mathcal{Y}=[C]$ for $T=C$ classes defining the possible labels. Let $\bm{t}(y)=(\llbracket y=1\rrbracket,\llbracket y=2\rrbracket,\ldots\llbracket y=C\rrbracket)$, where $\llbracket r\rrbracket=1$ when the predicate $r$ is true and $\llbracket r\rrbracket=0$ otherwise. Noting our results do not depend on minimal sufficiency, this $\bm{t}(y)$ is sufficient but not minimal sufficient. In this setting, the neural network outputs $\bm{h}$ correspond to the logits of the label probabilities. The resulting probabilities $p(y\,|\,\bm{x})$ is the softmax values of $\bm{h}$ denoted by $\sigma(\bm{x})\vcentcolon=\mathrm{SoftMax}(\bm{h}(\bm{x}))\in[0,1]^{T}$.
@@
Under this setting, we have $\mathcal{I}(\bm{h}\,|\,\bm{x})=\operatorname*{\mathrm{diag}}(\sigma(\bm{x}))-\sigma(\bm{x})\sigma(\bm{x})^{\top}$, whose eigenvalues do not follow a convenient pattern as the number of classes $C$ increases (Withers & Nadarajah, 2014). Likewise, the general structure of $\mathcal{K}(\bm{t}\,|\,\bm{x})$ does not provide a convinent equation for the maximum tensor eigenvalue. As such, we provide upper bounds for the maximum eigenvalues of $\mathcal{I}(\bm{h}\,|\,\bm{x})$ and $\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})-\mathcal{I}(\bm{h}\,|\,\bm{x})\otimes\mathcal{I}(\bm{h}\,|\,\bm{x})$ using Corollaries 4.2 and 4.3.

### v3

@@ 4 Practical Variance Estimation
$\bm{h}(\bm{x})=\bm{w}_{-1}^{j}\phi(\bm{h}^{\top}_{-2}(\bm{x})\bm{w}_{-2}^{j}+C_{-2})+\bm{c}_{-1}$ defines the NN mapping w.r.t. the $j$’th neuron in the second last layer, where $\bm{w}_{-2}^{j}$ and $\bm{w}_{-1}^{j}$ are incoming and outgoing links of the interested neuron, respectively; $\bm{h}_{-2}(\bm{x})$ is the output of the third last layer; and $\phi$ is the activation function. The ‘constants’ $C_{-2}$ and $\bm{c}_{-1}$ denote an aggregation of all terms which are independent of $\bm{w}_{-2}^{j}$ and $\bm{w}_{-1}^{j}$ in their respective layers. The Hessian of $\bm{h}_{k}(\bm{x})$ w.r.t. $\bm{w}_{-2}^{j}$ is $\partial^{2}\bm{h}_{k}(\bm{x})=(\bm{w}_{-1}^{j})_{k}\cdot\phi^{\prime\prime}(\bm{h}^{\top}_{-2}(\bm{x})\bm{w}_{-2}^{j}+C_{-2})\cdot(\bm{h}_{-2}(\bm{x})\bm{h}_{-2}^{\top}(\bm{x}))$. By Theorem 4.1, ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ can be arbitrarily small depending on $\phi^{\prime\prime}(\bm{h}^{\top}_{-2}(\bm{x})\bm{w}_{-2}^{j}+C_{-2})$. For example, if $\phi(t)=1/(1+\exp(-t))$, then $\phi^{\prime\prime}(t)=\phi(t)(1-\phi(t))(1-2\phi(t))$. In this case, for a neuron in the second last layer, a sufficient condition for ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})=0$ (and having ^$\hat{\mathcal{I}}_{2}$ favored against ^$\hat{\mathcal{I}}_{1}$) is $\bm{h}^{\top}_{-2}(\bm{x})\bm{w}_{-2}^{j}+C_{-2}=0$ for the neuron’s pre-activation. When the pre-activation value is saturated ($\infty$- ∞ or $\infty$), we also have that ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})={\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})=0$. Alternatively, suppose that $\phi(t)=\textrm{SoftPlus}(t)\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\log(1+\exp(t))$, a continuous relaxation of $\mathrm{ReLU}$, then $\phi^{\prime\prime}(t)=\phi^{\prime}(t)(1-\phi^{\prime}(t))$ where $\phi^{\prime}(t)=1/(1+\exp(-t))$. Then a sufficient condition for ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})=0$ with ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})\neq 0$ for a neuron in the second last layer is $\bm{h}^{\top}_{-2}(\bm{x})\bm{w}_{-2}^{j}+C_{-2}\rightarrow+\infty$.
These observations are further clarified by looking at related quantities over multiple parameters. So far we have only examined the variance of the FIM element-wise w.r.t. parameters $\theta_{i}$. To study all parameters $\bm{\theta}$ jointly, we consider the trace variances of the FIM estimators: for any $j\in\{1,2\}$, $\mathcal{V}_{j}(\bm{\theta}\,|\,\bm{x})$ denotes the trace of the covariance matrix of $\operatorname*{\mathrm{diag}}(\hat{\mathcal{I}}_{j}(\bm{\theta}\,|\,\bm{x}))$, where $\operatorname*{\mathrm{diag}}(\cdot)$ extracts a matrix’s diagonal elements into a column vector. We present upper bounds of these joint quantities.
@@
where $\mathrm{dHes}(\bm{h}\,|\,\bm{x})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(\operatorname*{\mathrm{diag}}(\mathrm{Hes}(\bm{h}_{1}\,|\,\bm{x})),\ldots,\operatorname*{\mathrm{diag}}(\mathrm{Hes}(\bm{h}_{T}\,|\,\bm{x})))$ and $\|\cdot\|_{F}$ is the Frobenius norm.
This upper bound comes from integrating the parameter-wise variances in Theorem 4.1 and incorporating a trace variance bound which utilizes the full spectrum of the NN derivatives and sufficient statistics quantities. This is fully depicted in Theorem D.1. Lower bounds can also be derived in terms of singular values (deferred to the Appendix). Note the upper bounds in Corollary 4.6 can be improved by expressing the $\min$ function’s first term with singular value quantities.
Having the $\min$ function in Corollary 4.6 is helpful as it shows a trade-off between two upper bounds: the scale of NN derivatives $\partial\bm{h}(\bm{x})$ and $\partial^{2}\bm{h}(\bm{x})$ versus the spectrum of the sufficient statistic terms. In the case of Eqs. 13 and 15, the trace of $\mathcal{I}(\bm{h}\,|\,\bm{x})$ is exactly the sum of all eigenvalues, including $\lambda_{\max}(\mathcal{I}(\bm{h}\,|\,\bm{x}))$. This can be helpful when the scale of the NN derivatives are not bounded by a small value. It should be noted that, by the chain rule, these NN derivatives scale with the overall sharpness / flatness (Li et al., 2018) of the landscape of the loss, i.e., the log-likelihood of Eq. 1. For NNs with large derivatives, the first term of the $\min$ could yield tight bounds of the variance, and one can therefore avoid dealing with the quadratic scaling of $\|\partial\bm{h}(\bm{x})\|$ in the second term. On the other hand, if the sharpness of the NN $\bm{h}$ can be controlled, e.g. via sharpness aware minimization (Foret et al., 2021), then one can benefit from the second term of the $\min$ and avoid computing the full spectrum of $\mathcal{I}(\bm{h}\,|\,\bm{x})$ in the first term.
Joint FIM Estimators
In the above, we considered the variance of conditional FIMs, which can scale differently depending on the input $\bm{x}$. Prior work’s analysis was limited to that of conditional FIMs (and their estimators) (Soen and Sun, 2021). Nevertheless, the ‘joint FIM’ estimators ^$\hat{\mathcal{I}}_{j}(\theta_{i})$ depend on sampling of $\bm{x}$ w.r.t. the data distribution $q(\bm{x})$. The bounds in Eq. 7 can be extended to the joint FIM $\mathcal{I}(\theta_{i})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\mathop{\mathbb{E}}_{q(\bm{x})}\mathcal{I}(\theta_{i}\,|\,\bm{x})$ by simply taking an expectation $\mathop{\mathbb{E}}_{q(\bm{x})}$ over the bounds. To analyze the variances of the joint FIM estimators ${\mathcal{V}}_{j}(\theta_{i})$, we present the following theorem which connects the prior results established for ${\mathcal{V}}_{j}(\theta_{i}\,|\,\bm{x})$, e.g. Theorem 4.1, via the law of total variance.
@@
For any $j\in\{1,2\}$, given $N_{x}$ samples of $\bm{x}\sim q(\bm{x})$ and $N$ samples of $\bm{y}_{\,|\,\bm{x}}\sim p(\bm{y}\,|\,\bm{x})$ for each $\bm{x}$ sampled,
${\mathcal{V}}_{j}(\theta_{i})=\frac{1}{N_{x}}\cdot\mathrm{Var}\left(\mathcal{I}(\theta_{i}\,|\,\bm{x})\right)+\frac{1}{N_{x}}\cdot\mathop{\mathbb{E}}_{q(\bm{x})}\left[{\mathcal{V}}_{j}(\theta_{i}\,|\,\bm{x})\right],$ (16)
where $\mathrm{Var}\left(\mathcal{I}(\theta_{i}\,|\,\bm{x})\right)$ is the variance of $\mathcal{I}(\theta_{i}\,|\,\bm{x})$ w.r.t. $q(\bm{x})$.
The dependence on $N$, the number of samples of ${\bm{y}}$ for each fixed ${\bm{x}}$, is hidden in ${\mathcal{V}}_{j}(\theta_{i}\,|\,\bm{x})$. When $N=1$, the hierarchical sampling described in the Theorem corresponds to an i.i.d. sampling of the joint distribution $p(\bm{x},\bm{y})$.
The variance incurred when estimating the FIM has two components. The first term on the RHS of Eq. 16 characterizes the randomness of the FIM w.r.t. $q(\bm{x})$, i.e., the input randomness. It vanishes when the FIM is estimated by taking the expectation w.r.t. $q(\bm{x})$, or the number of samples $\bm{x}$ is large enough. The second term (although also depending on $N_{x}$) comes from the sampling of $\bm{y}_{\,|\,\bm{x}}$ according to $p(\bm{y}\,|\,\bm{x})$, i.e., the output randomness, which scales with the central moments of $\bm{t}(\bm{y})$. If the NN is trained so that $p(\bm{y}\,|\,\bm{x})$ tends to be deterministic, this term will disappear leaving the first term to dominate. Eq. 16 can be further generalized using the law of total covariance to extend prior work considering conditional FIM covariances Soen and Sun (2021) to joint FIM covariances. Theorem 4.7 connects the variance of assuming a fixed input $\bm{x}$ with multiple samples $\bm{y}_{k}$ with the variance of pairs of samples $(\bm{x}_{k},\bm{y}_{k})$. The ${\mathcal{V}}_{j}(\theta_{i}\,|\,\bm{x})$ bounds in this section can thus be applied to the corresponding joint variance ${\mathcal{V}}_{j}(\theta_{i})$ by using this theorem. This is straightforward and omitted.
The first variance term in Theorem 4.7 is difficult to compute in practice: it relies on how the closed-form FIM varies w.r.t. $q(\bm{x})$. As such, it is useful to bound the first term into computable quantities.
@@
This upper bound is very similar to the 4th central moment term $\tilde{\lambda}_{\max}(\mathcal{K}^{p}(\bm{t}\,|\,\bm{x}))$ when considering the variance upper bound ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ in Theorems 4.1 and 4.2. In general, the eigenvalue terms of Lemma 4.8 and Theorem 4.1 are distinct, i.e., $\lambda^{2}_{\max}(\mathcal{I}(\bm{h}\,|\,\bm{x}))\neq\tilde{\lambda}_{\max}(\mathcal{K}^{p}(\bm{t}\,|\,\bm{x}))$. This is especially true for the classification and regression problems explored in this paper (see Table 1). However, the maximum eigenvalues can be related for exponential families with bounded sufficient statistic via Proposition 4.3, making both bounds depend only on $\lambda_{\max}(\mathcal{I}(\bm{h}\,|\,\bm{x}))$.
The total number of samples of $(\bm{x}_{k},\bm{y}_{k})$ is $N_{x}\cdot N$. In terms of sample complexity of the joint variance, using Theorems 4.7 and 4.8, the bound’s rate is given by $\mathcal{O}(1/N_{x}+1/(N_{x}\cdot N))$.
@@ 5 Case Studies
To make our theoretic results more concrete, we consider regression and classification settings, which correspond to specifying the exponential family in Eq. 1 to an isotropic Gaussian distribution and a categorical distribution, respectively. We include an empirical analysis of NNs trained on MNIST. Notably, our analysis considers general multi-dimensional NN output. This extends the case studies of Soen and Sun (2021) which was limited to 1D distributions due to the limitations of their bounds (and their associated computational costs of dealing with a 4D tensor of the full covariance).
Regression: Isotropic Gaussian Distribution
To characterize regression, we consider Gaussian distributions. As per Eq. 1, we have $\bm{t}(\bm{y})=\bm{y}\in\Re^{T}$ and base measure $\pi(\bm{y})\propto\exp(-\frac{1}{2}{\bm{y}}^{\top}{\bm{y}})$. This corresponds to the case where $\bm{h}(\bm{x})$ is trained via the squared loss. In this case, $\mathcal{I}(\bm{h}\,|\,\bm{x})=\bm{I}$ and $\mathcal{K}_{abcd}^{p}(\bm{t}\,|\,\bm{x})=\mathcal{I}_{ab}(\bm{h}\,|\,\bm{x})\cdot\mathcal{I}_{cd}(\bm{h}\,|\,\bm{x})+\,\mathcal{I}_{ac}(\bm{h}\,|\,\bm{x})\cdot\mathcal{I}_{bd}(\bm{h}\,|\,\bm{x})+\mathcal{I}_{ad}(\bm{h}\,|\,\bm{x})\cdot\mathcal{I}_{bc}(\bm{h}\,|\,\bm{x})$. The derivatives of the log-partition function $F(\bm{h})$ yields these central moments (Soen and Sun, 2021, Lemma 5). To apply Theorem 4.1, we examine the extreme eigenvalues of $\mathcal{I}(\bm{h}\,|\,\bm{x})$ and $\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})-\mathcal{I}(\bm{h}\,|\,\bm{x})\otimes\mathcal{I}(\bm{h}\,|\,\bm{x})$.
@@
Suppose that Eq. 1 is an isotropic Gaussian distribution. Then:
@@
$\tilde{\lambda}_{\min}(\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})-\mathcal{I}(\bm{h}\,|\,\bm{x})\otimes\mathcal{I}(\bm{h}\,|\,\bm{x}))=\tilde{\lambda}_{\max}(\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})-\mathcal{I}(\bm{h}\,|\,\bm{x})\otimes\mathcal{I}(\bm{h}\,|\,\bm{x}))={2}.$
@@
Hence, for regression the eigenvalues of sufficient statistics quantities in our bounds are equal. As such, in this case, the bound for $\mathcal{I}(\theta_{i}\,|\,\bm{x})$, ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$, and ${\mathcal{V}}_{2}(\theta_{i}\,|\,\bm{x})$ in Theorem 4.1 are all equalities.
@@
As $\mathcal{I}(\theta_{i}\,|\,\bm{x})$ can be written exactly in terms of the gradients of $\bm{h}$, in practice one does not need to utilize a random estimator when computing the conditional FIM, which simplifies to a Gauss-Newton matrix for the squared loss (Martens, 2020). When computing the FIM over a random sample $\bm{x}$, a variance still appears due to Theorem 4.7. By Lemma 4.8 and Proposition 5.1, the variance over a joint distribution is bounded by a function of the derivative $\partial_{i}\bm{h}(\bm{x})$ over the marginal input distributions: ${\mathcal{V}}(\theta_{i})\leq\mathop{\mathbb{E}}_{q(\bm{x})}[\|\partial_{i}\bm{h}(\bm{x})\|_{2}^{4}]\leq\max_{\bm{x}\in\mathrm{Supp}(q)}\|\partial_{i}\bm{h}(\bm{x})\|_{2}^{4}$. In other words, the overall variance to approximate the joint FIM is bounded by the gradients of the NN outputs.
Classification: Categorical Distribution
@@
For multi-class classification, we instantiate our exponential family with a categorical distribution (with $\pi(y)=1$). This corresponds to training a classifier NN with the log-loss. Let $\mathcal{Y}=[C]$ for $T=C$ classes defining the possible labels. Let $\bm{t}(y)=(\llbracket y=1\rrbracket,\ldots\llbracket y=C\rrbracket)$, where $\llbracket r\rrbracket=1$ when the predicate $r$ is true and $\llbracket r\rrbracket=0$ otherwise. Noting our results do not depend on minimal sufficiency, this $\bm{t}(y)$ is sufficient but not minimal sufficient. In this setting, the NN outputs $\bm{h}$ correspond to the logits of the label probabilities. The resulting probabilities $p(y\,|\,\bm{x})$ are the softmax values of $\bm{h}$ denoted by $\sigma(\bm{x})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\mathrm{SoftMax}(\bm{h}(\bm{x}))\in[0,1]^{T}$.
@@
Under this setting, we have $\mathcal{I}(\bm{h}\,|\,\bm{x})=\operatorname*{\mathrm{Diag}}(\sigma(\bm{x}))-\sigma(\bm{x})\sigma(\bm{x})^{\top}$ (where $\operatorname*{\mathrm{Diag}}(\sigma(\bm{x}))$ is the diagonal matrix with its diagonal entries set to $\sigma(\bm{x})$), whose eigenvalues do not follow a convenient pattern as $C$ increases (Withers and Nadarajah, 2014). Likewise, the maximum eigenvalue of $\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})$ is not available in simple closed form. As such, we provide upper bounds for the maximum eigenvalues of $\mathcal{I}(\bm{h}\,|\,\bm{x})$ and $\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})-\mathcal{I}(\bm{h}\,|\,\bm{x})\otimes\mathcal{I}(\bm{h}\,|\,\bm{x})$ using Corollaries 4.2 and 4.3.